The monthly cost (in dollars) of a long-distance phone plan is a linear function of the total calling time (in minutes). The monthly cost for 37
minutes of calls is $15.42
and the monthly cost for 53
minutes is $17.18
. What is the monthly cost for 39
minutes of calls?


Found 2 solutions by Theo, greenestamps:
Answer by Theo(13342)

linear function is y = mx + b in slope intercept form.
m is the slope
b is the y-intercept
the slope is the change in y divided by the corresponding change in x.
the y-intercept is the value of y when x = 0.

37 minutes = $15.42
53 minutes = $17.18

x is the minutes
y is the cost.

you have two points to consider.
they are (x1,y1) = (37,15.42) and (x2,y2) = 53,17.18)

the change in y is 17.18 - 15.42 = 1.76.
the corresponding change in x is 53 - 37 = 16.
the slope is 1.76 . 16 = .11

y = mx + b becomes y = .11 * x + b
replace x and y with either one of the coordinate points to solve for b.
i chose (37,15.42).
the equation becomes 15.42 = .11 * 37 + b.
solve for b to get b = 15.42 - .11 * 37 = 15.42 - 11.35

your equation becomes y = .11 * x + 11.35

when x = 0, y = 11.35
when x = 37, y = .11 * 37 + 11.35 = 15.42
when x = 53, y = .11 * 53 + 11.35 = 17.18

the equation looks good.
for 39 minutes, the total cost will be .11 * 39 + 11.35 = $15.64.

the equation can be graphed.
the equation to be graphed is y = .11 * x + 11.35
here's what the graph looks like.

$$$


Answer by greenestamps(13200)



The additional 53-37 = 16 minutes cost an additional $17.18-$15.42 = $1.76; the rate is $1.76/16 = $0.11 per minute.

The cost of 39 minutes of calls is the cost of 37 minutes of calls plus the cost of 2 more minutes at the rate of $0.11 per minute:

$15.42 + 2($0.11) = $15.42 + $0.22 = $15.64